### v4

@@ 2.1.3 Theoretical guarantees
The sole requirement for RPCholesky’s success is an eigenvalue decay condition which ensures the accuracy of the low-rank approximation $\bm{\widehat{A}}\approx\bm{A}$. To explain this condition in detail, we introduce a quantitative measure of the eigenvalue decay called the $\mu$-tail rank of the matrix $\bm{A}$:
@@
$\textup{rank}_{\mu}(\bm{A}):=\min\Bigl{\{}r\geq 0:\,\sum\nolimits_{i>r}\lambda_{i}(\bm{A})\leq\mu\Bigr{\}}.$
@@
Next, we state our main performance guarantee for RPCholesky preconditioning; the proof appears in Section 5.1.
@@
Fix a failure probability $\delta\in(0,1)$ and an error tolerance $\varepsilon\in(0,1)$. Let $\bm{A}$ be any positive-semidefinite matrix. Construct a random approximation $\bm{\widehat{A}}$ using RPCholesky with block size $B=1$ and approximation rank
@@
With probability at least $1-\delta$ over the randomness in $\bm{\widehat{A}}$, the RPCholesky preconditioner $\bm{P}=\bm{\widehat{A}}+\mu\mathbf{I}$ controls the condition number at a level
@@
Conditional on the event (2.5), when we apply preconditioned CG to the KRR linear system $(\bm{A}+\mu\mathbf{I})\bm{\beta}=\bm{y}$, we obtain an approximate solution $\bm{\beta}^{(t)}$ for which
$\lVert\bm{\beta}^{(t)}-\bm{\beta}\rVert_{\bm{A}+\mu\mathbf{I}}\leq\varepsilon\cdot\lVert\bm{\beta}\rVert_{\bm{A}+\mu\mathbf{I}}$ (2.6)
at any iteration $t\geq\delta^{-1/2}\log(2/\varepsilon)$, where $\bm{\beta}$ is the actual solution.
@@
Theorem 2.2 ensures that RPCholesky-preconditioned CG can solve any full-data KRR problem to fixed error tolerance $\varepsilon$ with failure probability $\delta$, provided that the eigenvalue condition (2.4) is satisfied. This eigenvalue condition ensures that the approximation rank $r$ is large enough to reliably capture the large eigenvalues of the kernel matrix. The factor $\log(\operatorname{tr}\bm{A}/\mu)$ is modest in size, being at most $\log\bigl{(}2^{53}\bigr{)}=36.7$ in double-precision arithmetic. Thus, the expression (2.4) is mainly determined by the $\textup{rank}_{\mu}(\bm{A})$ factor, which counts the number of “large” eigenvalues with the rest of the eigenvalues adding up to size $\mu$ or smaller.
@@
If $\textup{rank}_{\mu}(\bm{A})$ is $\mathcal{O}(\sqrt{N})$ and we apply RPCholesky preconditioning with $r=\mathcal{O}(\sqrt{N})$ columns, Theorem 2.2 guarantees that we can solve any full-data KRR problem in $\mathcal{O}(N^{2})$ operations. Uniform and greedy Nyström preconditioning admit no similar guarantee (see Section 3.1). Theorem 2.2 also provides insight into the case where the $\mu$-tail rank is larger than $\mathcal{O}(\sqrt{N})$. In this case, we may need to use a larger approximation rank $r$, and the construction cost for the preconditioner would exceed $\mathcal{O}(N^{2})$ operations.
Figure 3 evaluates RPCholesky’s capacity to control the conditioning of the 20 example problems when $\mu/N=10^{-7}$. Before RPCholesky preconditioning, the condition numbers range from $10^{4}$$10^{7}$. RPCholesky typically reduces the condition number to $10^{1}$$10^{3}$ and sometimes even to $10^{0}$. However, RPCholesky cannot fully control the conditioning for the most difficult problems, as no rank-$r$ Nyström preconditioner can reduce the condition number beyond $(\lambda_{r+1}(\bm{A})+\mu)/(\lambda_{N}(\bm{A})+\mu)$. To further reduce the condition number, it would be necessary to significantly raise the approximation rank.
@@
Figure 3: Eigenvalue decay for the 20 KRR problems in Table 1. Left panel shows eigenvalues of the matrix $\bm{A}+\mu\mathbf{I}$ before RPCholesky preconditioning. Right panel shows eigenvalues of the matrix $\mu\bm{P}^{-1/2}(\bm{A}+\mu\mathbf{I})\bm{P}^{-1/2}$ after RPCholesky preconditioning with $r=1000$. Dashed line indicates the regularization parameter $\mu=10^{-7}N$.
@@
In practice, the parameter $r$ should be tuned to balance the cost of forming the preconditioner and the cost of the CG iterations. As a simple default, we recommend setting $r=10\sqrt{N}$. This default is 1.2×1.2$\times$1.2 × larger than the approximation rank used in Figure 3 and is large enough to ensure that all 20 test problems are solved in fewer than 200 iterations when $\mu/N=10^{-7}$. See Section 4.1 for more exploration of the parameter $r$ with a scientific data set.
@@ 2.2 Restricted kernel ridge regression
$\widehat{f}(\bm{x};\bm{\widehat{\beta}})=\sum\nolimits_{i=1}^{k}\widehat{\beta}_{i}K(\bm{x}^{(s_{i})},\bm{x}),$
using a subset $\bm{x}^{(s_{1})},\bm{x}^{(s_{2})},\ldots,\bm{x}^{(s_{k})}$ of input points, which are called “centers”. There are many strategies for selecting the centers, such as uniform sampling, ridge leverage score sampling, and RPCholesky. One must balance the computational cost of the center selection procedure against the quality of the centers. For all experiments in this paper, we use the computationally trivial approach of sampling centers uniformly at random; however, the development of fast procedures for identifying high-quality centers is an interesting topic for future work.
@@
Equation 2.7 is a $k\times k$ linear system that is smaller than the system eq. 2.1 for the full-data KRR problem. This system can become highly ill-conditioned. To mitigate this issue, we add a small multiple of the machine precision $\varepsilon_{\rm mach}$ to the diagonal entries of the regularizer $\mu\bm{A}(\mathsf{S},\mathsf{S})$.
Solving (2.7) by direct methods is expensive because of the $\mathcal{O}(k^{2}N)$ cost of forming the Gram matrix $\bm{G}=\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})$. When the number of centers is moderately large ($k\geq 10^{3}$), we recommend a cheaper approach for solving (2.7) using conjugate gradient with the KRILL preconditioner, described below. We will prove that KRILL preconditioning solves every restricted KRR problem in $\mathcal{O}((N+k^{2})k\log k)$ arithmetic operations (Theorem 2.3).
@@ 2.2.1 KRILL preconditioning
1:Positive semidefinite matrix $\bm{A}\in{}^{N\times N}$, right-hand-side vector $\bm{y}\in{}^{N}$, regularization coefficient $\mu$, centers $\mathsf{S}$, and tolerance $\varepsilon$.
@@
3:$\bm{H}\leftarrow\mu\bm{A}(\mathsf{S},\mathsf{S})$
4:$\bm{H}\leftarrow\bm{H}+N\varepsilon_{\rm mach}\operatorname{tr}(\bm{A}(\mathsf{S},\mathsf{S}))\cdot\mathbf{I}$. $\triangleright$ $\varepsilon_{\rm mach}\approx 2\times 10^{-16}$ in double precision.
5:$k\leftarrow|\mathsf{S}|$.
6:$\bm{\Phi}\leftarrow\textsc{SparseSignEmbedding}(d=2k,N,\zeta=\min\{8,2k\})$. $\triangleright$ See Algorithm 5.
7:$\bm{B}\leftarrow\bm{\Phi}\bm{A}(:,\mathsf{S})$.
8:$\bm{P}\leftarrow\bm{B}^{\ast}\bm{B}+\bm{H}$.
9:$\bm{C}\leftarrow\textsc{Cholesky}(\bm{P}+\varepsilon_{\rm mach}\operatorname{tr}(\bm{P})\cdot\mathbf{I})$.
@@
Preconditioner $:\widehat{\bm{\beta}}\mapsto\bm{C}^{-1}(\bm{C}^{-\ast}\widehat{\bm{\beta}}).$
11:$\bm{\widehat{\beta}}_{\star}\leftarrow\textsc{PCG}(\texttt{Product},\bm{A}(\mathsf{S},:)\bm{y},\varepsilon,\texttt{Preconditioner})$. $\triangleright$ See Algorithm 3.
The KRILL preconditioner is based on a randomized approximation of the Gram matrix $\bm{G}=\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})$. To form this approximation, first generate a sparse random sign embedding [29, §9.2]:
$\bm{\Phi}=\tfrac{1}{\sqrt{\zeta}}\begin{bmatrix}\bm{\phi}_{1}&\cdots&\bm{\phi}_{N}\end{bmatrix}\in{}^{d\times N},$
where $\bm{\phi}_{1},\ldots,\bm{\phi}_{N}$ are sparse columns that possess uniform ±$\pm 1$ values in $\zeta$ uniformly randomly chosen positions. Then, form the matrix $\bm{B}=\bm{\Phi}\bm{A}(:,\mathsf{S})$ and compute the outer product matrix $\bm{\widehat{G}}=\bm{B}^{*}\bm{B}$. Last, add a small multiple of $\bm{A}(\mathsf{S},\mathsf{S})$ to $\bm{\widehat{G}}$ to construct the KRILL preconditioner:
@@
The pseudocode for KRILL preconditioning is provided in Algorithm 2.
KRILL preconditioning is efficient because evaluating the approximate Gram matrix $\bm{\widehat{G}}$ requires fewer operations than evaluating the exact Gram matrix $\bm{G}$. The precise number of operations depends on the sparsity $\zeta$ and the embedding dimension $d$. In our theoretical analysis, we assume $\zeta=\mathcal{O}(\log k)$ and $d=\mathcal{O}(k\log k)$, and we can form $\bm{\widehat{G}}$ in $\mathcal{O}((N+k^{2})k\log k)$ operations. In practice, we still observe fast CG convergence when we use a smaller, sparser embedding with $\zeta=\min\{8,2k\}$ and $d=2k$ [50, p. A2440]. Consequently, after including the cost of the CG iterations, the total operating cost of KRILL preconditioning is $\mathcal{O}((N+k^{2})k\log k)$ in theory and $\mathcal{O}((N+k^{2})k)$ in practice.
@@ 2.2.2 Empirical performance
Figure 4 evaluates the performance of KRILL across the 20 regression and classification problems described in Table 1. For each problem, we subsample $N=4\times 10^{4}$ data points and select $k=1000$ centers uniformly at random. We then standardize the features and apply KRR using the squared exponential kernel (2.3) with the bandwidth $\sigma=3$. The regularization is either set to a moderate value $\mu/N=10^{-6}$ or a tiny value $\mu/N=10^{-12}$. We run $100$ CG iterations and declare a problem to be “solved” if the relative residual
@@
falls below a tolerance of $\varepsilon=10^{-4}$. For comparison, Figure 4 also evaluates the performance of the current state-of-the-art solver FALKON, which approximates the Gram matrix $\bm{G}$ via Monte Carlo sampling. For essential details of FALKON, see Section 3.2.
Moderate regularization: $\mu=10^{-6}\cdot N$
Tiny regularization: $\mu=10^{-12}\cdot N$
Figure 4: Fraction of solved problems versus number of CG iterations for the 20 kernel problems in Table 1.
Examining Figure 4, we find that KRILL solves all of the classification and regression problems in 30 iterations, both when the regularization is moderate (left) and when the regularization is tiny (right). In contrast, FALKON struggles with the tiny regularization (right panel), and it solves just 9 problems after 30 CG iterations.
KRILL is a randomized algorithm, so it might change behavior depending on the random seed. Yet empirically, KRILL produces consistent results up to ±$\pm 10\%$ random variation in the number of CG iterations. Figure 5 shows the median and the 20%–80% quantiles of the relative residual eq. 2.9 for the problems with the fastest (yolanda, left) and slowest (creditcard, right) CG convergence. The quantiles are calculated using 100 random subsamples $\mathsf{S}$ and subsequent CG solves. In these problems, the 20%–80% quantiles are highly concentrated around the median.
yolanda Refer to caption
creditcard Refer to caption
Figure 5: Relative residual versus number of CG iterations for the problems with the fastest (yolanda, left) and slowest (creditcard, right) CG convergence when $\mu/N=10^{-6}$. Thick lines show the median error and shaded regions show the $20\%$$80\%$ error quantiles over 100 random trials.

### v5

@@ 2.1.3 Theoretical guarantees
RPCholesky is guaranteed to solve any full-data KRR problem given sufficiently fast eigenvalue decay in the kernel matrix $\bm{A}$. To explain the eigenvalue decay condition in detail, we introduce a quantitative measure called the $\mu$-tail rank:
@@
$\textup{rank}_{\mu}(\bm{A}):=\min\Biggl\{r\geq 0:\,\sum_{i>r}\lambda_{i}(\bm{A})\leq\mu\Biggr\}.$
@@
Next, we state our main performance guarantee for RPCholesky preconditioning; the proof appears in Section˜5.1.
@@
Fix a failure probability $\delta\in(0,1)$ and an error tolerance $\varepsilon\in(0,1)$. Let $\bm{A}$ be any positive-semidefinite matrix, and let $\mu>0$ be a positive number. Construct a random approximation $\bm{\widehat{A}}$ using RPCholesky with block size $B=1$ and approximation rank
@@
With probability at least $1-\delta$, the RPCholesky preconditioner $\bm{P}=\bm{\widehat{A}}+\mu\mathbf{I}$ controls the condition number at a level
@@
Conditional on the event (2.5), when we apply preconditioned CG to the KRR linear system $(\bm{A}+\mu\mathbf{I})\bm{\beta}=\bm{y}$, we obtain an approximation $\bm{\beta}^{(t)}$ to the actual solution $\bm{\beta}$ that satisfies
$\lVert\bm{\beta}^{(t)}-\bm{\beta}\rVert_{\bm{A}+\mu\mathbf{I}}\leq\varepsilon\,\lVert\bm{\beta}\rVert_{\bm{A}+\mu\mathbf{I}}$ (2.6)
[…]
@@
Theorem˜2.2 ensures that RPCholesky-preconditioned CG can solve any full-data KRR problem up to a fixed error tolerance $\varepsilon$ with failure probability $\delta$ when the eigenvalue condition (2.4) is satisfied. This eigenvalue condition ensures that the approximation rank $r$ is large enough to reliably capture the large eigenvalues of the kernel matrix. The factor $\log(\operatorname{tr}\bm{A}/\mu)$ is modest in size, so the expression (2.4) is mainly determined by $\textup{rank}_{\mu}(\bm{A})$, which counts the number of “large” eigenvalues with the rest of the eigenvalues adding up to size $\mu$ or smaller.
Theorem˜2.2 assumes RPCholesky is performed with a block size $B=1$ for simplicity. This theorem can be extended to $B>1$ given a slightly higher approximation rank:
$r\geq\bigl(\textup{rank}_{\mu}(\bm{A})+B\bigr)\,\bigl(1+\log\bigl(\operatorname{tr}\bm{A}/\mu\bigr)\bigr).$
The recent work [24] analyzes RPCholesky with arbitrary block size $B\geq 1$, and it introduces an “accelerated RPCholesky” algorithm that improves the accuracy of block RPCholesky.
@@
If the $\mu$-tail rank is $\mathcal{O}(\sqrt{N})$ and we apply RPCholesky preconditioning with $\mathcal{O}(\sqrt{N})$ columns, Theorem˜2.2 guarantees that we can solve any full-data KRR problem in $\mathcal{O}(N^{2})$ operations. Uniform and greedy Nyström preconditioning admit no similar guarantee (see Section˜3.1). Theorem˜2.2 also provides insight into the case where the $\mu$-tail rank is larger than $\mathcal{O}(\sqrt{N})$. In this case, we may need to use a larger approximation rank $r$, and the construction cost for the preconditioner would exceed $\mathcal{O}(N^{2})$ operations.
Figure˜3 evaluates RPCholesky’s impact on the 20 example problems when $\mu/N=10^{-7}$. Before RPCholesky preconditioning, the condition number $\lambda_{\max}/\lambda_{\min}$ is nearly $10^{7}$ (left panel). RPCholesky reduces the condition number to a level $10^{0}$$10^{4}$ (middle panel). The impact of RPCholesky is similar to the optimal rank-$r$ preconditioner, which eliminates the leading $r$ eigenvalues by setting them equal to $\mu$ (right panel). Yet even the optimal rank-$r$ preconditioner is sensitive to slow kernel matrix eigenvalue decay.
No preconditioner       RPCholesky eigenvalues     Eigenvalues shifted
eigenvalues $\lambda_{i}(\bm{A}+\mu\mathbf{I})$    $\lambda_{i}(\mu\bm{P}^{-1/2}(\bm{A}+\mu\mathbf{I})\bm{P}^{-1/2})$    by $r$ indices $\lambda_{i+r}(\bm{A}+\mu\mathbf{I})$
@@
Figure 3: Eigenvalue decay for the 20 KRR problems in Table˜1. Left panel shows eigenvalues of the matrix $\bm{A}+\mu\mathbf{I}$ before RPCholesky preconditioning. Middle panel shows eigenvalues of the matrix $\mu\bm{P}^{-1/2}(\bm{A}+\mu\mathbf{I})\bm{P}^{-1/2}$ after RPCholesky preconditioning with $r=1000$. Right panel shows eigenvalues $\lambda_{i+r}(\bm{A}+\mu\mathbf{I})$ resulting from the mathematically optimal rank-$r$ preconditioner. Dashed lines indicate the regularization parameter $\mu=10^{-7}N$.
@@
The best application of RPCholesky balances the cost of forming the preconditioner with the cost of the resulting preconditioned CG iterations. As a simple default, we recommend running RPCholesky with $r=10\sqrt{N}$. This default is $1.2\times$ larger than the approximation rank used in Figure˜3 and is large enough to ensure that all 20 test problems are solved in fewer than 120 iterations when $\mu/N=10^{-7}$. See Section˜4.1 for more exploration of the parameter $r$ with a scientific data set.
@@ 2.2 Restricted kernel ridge regression
$\widehat{f}(\bm{x};\bm{\widehat{\beta}})=\sum_{i=1}^{k}\widehat{\beta}_{i}K(\bm{x}^{(s_{i})},\bm{x}),$
using a subset $\bm{x}^{(s_{1})},\bm{x}^{(s_{2})},\ldots,\bm{x}^{(s_{k})}$ of input points, which are called “centers”. There are many strategies for selecting the centers, such as uniform sampling, ridge leverage score sampling, and RPCholesky. One must balance the computational cost of the center selection procedure against the quality of the centers. For all the experiments in this paper, we use the computationally trivial approach of sampling centers uniformly at random. However, the development of fast procedures for identifying high-quality centers is a topic for future work.
@@
Equation (2.7) is a $k\times k$ linear system that is potentially much smaller than the $N\times N$ system (2.1) in full-data KRR. This system can be arbitrarily ill-conditioned and may even be singular. Therefore, to accurately solve restricted KRR problems in finite-precision arithmetic, we add a small multiple of the machine precision to the diagonal entries of the regularizer $\mu\bm{A}(\mathsf{S},\mathsf{S})$.
Solving (2.7) by direct methods is expensive because of the $\mathcal{O}(k^{2}N)$ cost of forming the Gram matrix $\bm{G}=\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})$. When the number of centers is $k\geq 10^{3}$, we recommend a cheaper approach for solving (2.7) using conjugate gradient with the KRILL preconditioner, described below. We will prove that KRILL preconditioning solves every restricted KRR problem in $\mathcal{O}((N+k^{2})\,k\log k)$ arithmetic operations (Theorem˜2.3).
@@ 2.2.1 KRILL preconditioning
1:Positive semidefinite matrix $\bm{A}\in\real^{N\times N}$, right-hand-side vector $\bm{y}\in\real^{N}$, regularization coefficient $\mu$, centers $\mathsf{S}$, and tolerance $\varepsilon$.
@@
[…]
@@
[…]
KRILL is based on a randomized approximation of the Gram matrix $\bm{G}=\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})$. To form this approximation, we first generate a sparse random sign embedding [33, §9.2]:
$\bm{\Phi}=\tfrac{1}{\sqrt{\zeta}}\begin{bmatrix}\bm{\phi}_{1}&\cdots&\bm{\phi}_{N}\end{bmatrix}\in\real^{d\times N},$
where $\bm{\phi}_{1},\ldots,\bm{\phi}_{N}$ are columns that possess uniform $\pm 1$ values in $\zeta$ uniformly random positions. Next, we construct the matrix $\bm{B}=\bm{\Phi}\bm{A}(:,\mathsf{S})$. Last, to form the KRILL preconditioner, we calculate the approximate Gram matrix $\bm{\widehat{G}}=\bm{B}^{*}\bm{B}$ and add the regularizer $\mu\bm{A}(\mathsf{S},\mathsf{S})$:
@@
The pseudocode for KRILL preconditioning is provided in Algorithm˜2.
KRILL preconditioning is efficient because evaluating the approximate Gram matrix $\bm{\widehat{G}}$ requires fewer operations than evaluating the exact Gram matrix $\bm{G}$. The precise number of operations depends on the sparsity $\zeta$ and the embedding dimension $d$. In our theoretical analysis, we assume $\zeta$ and $d$ are functions given in [16] that grow according to $\zeta=\mathcal{O}(\log k)$ and $d=\mathcal{O}(k\log k)$. In practice, we set $\zeta=\lceil\log(k+1)\rceil$ and $d=2k$, which we have found to be sufficient to obtain fast CG convergence on all instances. (Smaller values for the sparsity parameter, such as $\zeta=4$, seems to be reliable as well [9].) Recent research [54, 14] has improved the theoretical understanding of sparse random sign embeddings, but the optimal sparsity for a given embedding dimension remains uncertain. After including the cost of the CG iterations, the total operating cost of KRILL preconditioning is therefore $\mathcal{O}((N+k^{2})\,k\log k)$ in theory and $\mathcal{O}(Nk\log k+k^{3})$ in practice.
@@ 2.2.2 Empirical performance
Figure˜4 evaluates the performance of KRILL across the 20 regression and classification problems described in Table˜1. For each problem, we subsample $N=4\times 10^{4}$ data points and select $k=1000$ centers uniformly at random. We standardize the features and apply KRR using the squared exponential kernel (2.3) with the bandwidth $\sigma=\sqrt{d}$. We either set the regularization to a large value $\mu/N=10^{-6}$ or a tiny value $\mu/N=10^{-12}$. We run $100$ CG iterations and declare a problem to be “solved” if the relative residual
@@
falls below a tolerance of $\varepsilon=10^{-4}$. For comparison, Figure˜4 also evaluates the performance of the FALKON preconditioner that approximates the Gram matrix $\bm{G}$ via Monte Carlo sampling. More details about FALKON are in Section˜3.2.
Large regularization: $\mu=10^{-6}N$
Tiny regularization: $\mu=10^{-12}N$
Figure 4: Fraction of solved problems versus number of CG iterations for the 20 kernel problems in Table˜1.
Examining Figure˜4, we find that KRILL solves all classification and regression problems in 30 or fewer iterations, both when the regularization is large (left) and when the regularization is tiny (right). In contrast, FALKON struggles with the tiny regularization (right panel), and it solves just 8 problems after 30 CG iterations.
KRILL is a randomized algorithm, so its behavior depends on the random seed. Yet the results of KRILL are consistent up to $\pm 20\%$ random variation in the number of CG iterations required for convergence. Figure˜5 shows the median and the 20%–80% quantiles of the relative residual (2.9) for the problems with the fastest (COMET_MC_SAMPLE, left) and slowest (creditcard, right) CG convergence. The quantiles are calculated using 100 independent trials including random subsampling of the data set and random preconditioner formation. The quantiles are tightly concentrated around the median.
COMET_MC_SAMPLE
creditcard
Figure 5: Relative residual versus number of CG iterations for the problems with the fastest (COMET_MC_SAMPLE, left) and slowest (creditcard, right) CG convergence when $\mu/N=10^{-6}$. Thick lines show the median and shaded regions show the $20\%$$80\%$ error quantiles over 100 random trials.